OpenStudy (anonymous):

how to determine a & b valuesin the equation: y=ax^2 +bx-4 if the vertex is located at (-2,-1) Please help!

myininaya (myininaya):

did you enter in -2 for x and -1 for y

hero (hero):

4a = b

hero (hero):

-1 = a(-2)^2 +(4a)(-2) - 4

hero (hero):

a = -3/4?

hero (hero):

b = 3?

hero (hero):

How did you get b = 2?

myininaya (myininaya):

so a has different value too

hero (hero):

myininaya, how come you just can't do things the normal way?

OpenStudy (anonymous):

so you cant just sub -2 into one equation and -1 into another, andthe other number equal zero? or does that not mke any sense?

OpenStudy (anonymous):

I'm just a little confused by the math you did above..lol

hero (hero):

myininaya likes to go overboard with her methods

myininaya (myininaya):

ok hero but there are two possible outcomes

hero (hero):

Not if you do it the way I did. I used direct substitution

myininaya (myininaya):

We have a parabola concave up and also concave down at vertex

myininaya (myininaya):

Here I will do it the way I told you to

OpenStudy (anonymous):

Im familar with substitution , just not the way shown above im only in gr 10 so mabye thats why i dont get yours?

hero (hero):

a = -3/4 b = -3

OpenStudy (anonymous):

^did you make 2 equations and then substitutions?

hero (hero):

something like that

OpenStudy (anonymous):

-1 = a(-2)^2 +(4a)(-2) - 4 :this is what you did, why did you sub 4a into the equation? is there a more strait forward way to do this? :)

hero (hero):

myininaya, if you use the direct substitution method, you wouldn't have to worry about getting hung up like that. The parabola opens upward or downward depending on the value of a. Since I calculated a to be negative, I'm pretty sure it opens downward and there is only one possibility

hero (hero):

I replace b with 4a because b = 4a

hero (hero):

I also replaced x = -2 and y = -1

hero (hero):

Which are all of the known values

hero (hero):

according to x = -b/2a, there's only one possible a

OpenStudy (anonymous):

alright im still a bit confused but thanks for your help!! :)

myininaya (myininaya):

\[-1=a(-2)^2+b(-2)-4\] \[-1=4a-2b-4 => 3=4a-2b\] We also know the vertex which is (-2,-1) \[y=a(x-h)^2+k\] we also know h=-2 and k=-1 so we have \[y=a(x+2)^2-1=a(x^2+4x+4)-1=ax^2+4ax+4a-1=ax^2+4ax+(4a-1)\] \[=>4a=b , -4=4a-1\] \[=>3=b-2b=> 3=-b=> b=-3\] if b=-3 then a=-3/4 hmmm how do we get the other possibility ok so maybe there is only one lol

hero (hero):

:D

hero (hero):

There's only one possibility because there's only one a value that would work in order to get both sides equal for x = -b/2a

hero (hero):

if a was squared, then there would be two possibilities

myininaya (myininaya):

I didn't think about the y-intercept

myininaya (myininaya):

The parabola has vertex (-2,-1) and -4 is lower so this graph will be concave down

myininaya (myininaya):



hero (hero):

yup

myininaya (myininaya):

we can't not have it concave up is has to come to that -4

myininaya (myininaya):

so my bad hero you win this one

hero (hero):

I don't win many so I will savor this :P

myininaya (myininaya):

hey lady do you understand what i did up there?

OpenStudy (anonymous):

A bit lol but thanks!

hero (hero):

mya likes to confuse people then ask if they understand

myininaya (myininaya):

you also have to use the vertex form like i did to find another equation or two

OpenStudy (anonymous):

im still pretty confused but i can move on to a different qestion on my assignment lol thanks for the help on this one tho :)

hero (hero):

You're still here?